OpenStudy (anonymous):

x/3+x/4=21/12 solve for x?

OpenStudy (anonymous):

add first \[\frac{x}{3}+\frac{x}{4}=\frac{3x+4x}{3\times 4}=\frac{7x}{12}\]

OpenStudy (anonymous):

if \(\frac{7x}{12}=\frac{21}{12}\) then you know \(7x=21\) and you can find \(x\)

OpenStudy (compassionate):

\[\frac{ x }{ 3 } + \frac{x}{4} = \frac{21}{12}\] You want to cross multiply and add denominators.

OpenStudy (compassionate):

Satelite has the right idea.

OpenStudy (anonymous):

Lol I still don't understand it

OpenStudy (anonymous):

we can try it another way you want to get \(x\) by itself, but you have these annoying fractions, so lets get rid of them

OpenStudy (anonymous):

you are given \[\frac{ x }{ 3 } + \frac{x}{4} = \frac{21}{12}\] and the least common multiple of 3, 4 and 12 is 12, so you can get rid of the fractions by multiplying both sides of the equation by 12

OpenStudy (anonymous):

\[12\times \left(\frac{ x }{ 3 } + \frac{x}{4} \right)=12\times \frac{21}{12}\]

OpenStudy (anonymous):

the right hand side is \(21\) because you are cancelling the \(12\) the left hand side is \(4x+3x=7x\) because 3 goes in to 12 4 times and 4 goes in to 12 3 times

OpenStudy (anonymous):

that gives you \(7x=21\) and if \(7x=21\) it is pretty clear what \(x\) is , right?